OpenStudy (anonymous):

Please help! If r = 3 and the sum of the 6 terms of the geometric series is 3,640, what is the first term?

OpenStudy (curry):

25

OpenStudy (anonymous):

2.5? Or 25?

OpenStudy (curry):

25

OpenStudy (lalaly):

Formula for Geometric sum Sn = a( r^(n) -1)/(r-1) Sn = 3640 r=3 n=6 Hence 3640 = a(3^6 -1)/(3-1) 3640 = a( 729-1)/2 3640=a(728/2) 3640=a(364) a= 10 a= 3640/1093